Anna is 12 years old, and her mother is three times as old as Anna. How many years ago was her mother four times as old as Anna?
Answer by stanbon(75887)

Anna is 12 years old, and her mother is three times as old as Anna. How many years ago was her mother four times as old as Anna?
-----
Equations:
a = 12
m = 3a = 3*12 = 36
---------------------------
(m-x)= 4(a-x)
---
36-x = 4(12-x)
36-x = 48-4x
3x = 12
x = 4 (# of years ago)
============================
Cheers,
Stan H.